# Autoencoding Random Forests

### v1

@@ 2 Background
Our starting point is the well established connection between RFs and kernel methods [13, 24, 78]. The basic insight is that classification and regression trees (CART) [15], which serve as basis functions for many popular ensemble methods, are a kind of adaptive nearest neighbors algorithm [61]. At the root of the tree, all samples are connected. Each split severs the link between one subset of the data and another (i.e., samples routed to left vs. right child nodes), resulting in a gradually sparser graph as depth increases. At completion, a sample’s “neighbors” are just those datapoints that are routed to the same leaf. Given some feature space $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, the implicit kernel of tree $b$, $k^{(b)}:\mathcal{X}\times\mathcal{X}\mapsto\{0,1\}$, is an indicator function that evaluates to $1$ for all and only neighboring sample pairs. [^1] This ignores certain subtleties that arise when trees are grown on bootstrap samples, in which case $k^{(b)}$ may occasionally evaluate to larger integers. For present purposes, we presume that trees are grown on data subsamples; see Appx. A. This base kernel can be used to define different ensemble kernels. For instance, taking an average over $B>1$ trees, we get a kernel with a simple interpretation as the proportion of trees in which two samples colocate: $k^{S}(\mathbf{x},\mathbf{x}^{\prime})=B^{-1}\sum_{b=1}^{B}k^{(b)}(\mathbf{x},\mathbf{x}^{\prime})$. We call this the Scornet kernel after one of its noted proponents [78], who showed that $k^{S}$ is provably close in expectation to the Breiman RF kernel $k_{n}^{RF}$:
@@
where $i\in[n]:=\{1,\dots,n\}$ indexes the training samples. This kernel represents the average of normalized tree kernels, and fully encodes the information learned by the RF $f_{n}$ via the identity:
$f_{n}(\mathbf{x})=\sum_{i=1}^{n}k^{RF}_{n}(\mathbf{x},\mathbf{x}_{i})\leavevmode\nobreak\ y_{i},$ (2)
@@
Previous authors have explored feature engineering with RFs. Shi and Horvath [80] perform multidimensional scaling on a dissimilarity matrix extracted from supervised and unsupervised forests. However, they do not explore the connections between this approach and kernel methods, nor do they propose any strategy for decoding latent representations. Other more heuristic approaches involve running PCA on a weighted matrix of all forest nodes (not just leaves), a method that works well in some experiments but comes with no formal guarantees [71].
Several generative algorithms based on RFs have been proposed, building on the probabilistic circuit literature [23, 94]. These do not involve any explicit encoding step, although the models they train could be passed through our pipeline to extract implicit embeddings (see Sect. 5). Existing methods for sum-product network encoding could in principle be applied to an RF following compilation into a corresponding circuit [90]. However, these can often increase rather than decrease dimensionality, and do not come with any associated decoding procedure.
Perhaps the most similar method to ours, in motivation if not in execution, is Feng and Zhou [30]’s encoder forest (eForest). This algorithm maps each point to a hyperrectangle defined by the intersection of all leaves to which the sample is routed. Decoding is then achieved by taking some representative value for each feature in the subregion (e.g., the median). Notably, this approach does not include any dimensionality reduction. On the contrary, the embedding space requires minima and maxima for all input variables, resulting in a representation with double the number of features as the inputs. Working from the conjecture that optimal prediction is equivalent to optimal compression [73, 53, 44], we aim to represent the information learned by the RF in relatively few dimensions.
@@ 3 Encoding
As a preliminary motivation, we prove several important properties of the RF kernel. The following definitions are standard in the literature. Let $\mathcal{X}$ be a compact metric space, and let $C(\mathcal{X})$ be the set of all real-valued continuous functions on $\mathcal{X}$.
@@
A symmetric function $k:\mathcal{X}\times\mathcal{X}\mapsto\mathbb{R}$ is positive semidefinite (PSD) if, for all $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in\mathcal{X}$, $n\in\mathbb{N}$, and $c_{i},c_{j}\in\mathbb{R}$, we have: $\sum_{i,j}^{n}c_{i}\leavevmode\nobreak\ c_{j}\leavevmode\nobreak\ k(\mathbf{x}_{i},\mathbf{x}_{j})\geq 0$.
@@
[…]
@@
$\lim_{n\rightarrow\infty}\leavevmode\nobreak\ P\big{(}\lVert f^{*}-f_{n}\rVert_{\infty}\geq\epsilon\big{)}=0.$
@@
The literature on kernel methods is largely focused on fixed kernels such as the radial basis function (RBF). Among adaptive partitioning alternatives, the Scornet kernel been studied in some detail [24, 78, 70], as has the Mondrian kernel [59, 4, 17]. However, to the best of our knowledge, we are the first to establish these results for the RF kernel. Thm. 3.4 confirms that $k_{n}^{RF}$ is flexible, informative, and generally “well-behaved” in ways that will prove helpful for autoencoding.
@@
$\min_{\mathbf{V}\in\mathbb{R}^{n\times d_{\mathcal{Z}}}}\sum_{i,j}k_{ij}\leavevmode\nobreak\ \lVert\mathbf{v}_{i}-\mathbf{v}_{j}\rVert_{2}\quad\text{s.t.}\leavevmode\nobreak\ \mathbf{V}^{\top}\mathbf{V}=\mathbf{I},$
@@ 4 Decoding
$\lim_{n\rightarrow\infty}\leavevmode\nobreak\ P\Big{(}\lVert\mathbf{x}-h_{n}\big{(}g^{*}(\mathbf{x})\big{)}\rVert_{\infty}\geq\epsilon\Big{)}=0.$
@@
[…]
@@ 4.1 Constrained Optimization
Now we must identify the most likely leaf assignments for each of our $m$ (unseen) test samples $\mathbf{X}_{0}\in\mathbb{R}^{m\times d_{\mathcal{X}}}$, given their latent representation $\mathbf{Z}_{0}$. Call this target matrix $\bm{\Psi}\in\{0,1\}^{m\times d_{\Phi}}$. To estimate it, we start with the binary matrix of leaf assignments for training samples $\bm{\Pi}\in\{0,1\}^{n\times d_{\Phi}}$. Exploiting the inner product definition of a kernel, observe that our original (training) adjacency matrix $\mathbf{K}$ satisfies $B\mathbf{K}=\bm{\Phi}\bm{\Phi}^{\top}=\bm{\Pi}\mathbf{S}\bm{\Pi}^{\top}$, where $\bm{\Phi}\in[0,1]^{n\times d_{\Phi}}$ is the RKHS representation of $\mathbf{X}$, and $\mathbf{S}=\text{diag}(\mathbf{s})$. We partition $[d_{\Phi}]$ into $B$ subsets $L^{(b)}$ that index the leaves belonging to tree $b$. Recall that each tree $b$ partitions the feature space $\mathcal{X}$ into $L^{(b)}$ hyperrectangular subregions $\mathcal{X}_{\ell}^{(b)}\subset\mathcal{X}$, one for each leaf $\ell\in[L^{(b)}]$. Let $R_{i}^{(b)}\in\{\mathcal{X}_{1}^{(b)},\dots,\mathcal{X}_{L^{(b)}}^{(b)}\}$ denote the region to which sample $i$ is routed in tree $b$. Then leaf assignments for test samples can be calculated by solving the following integer linear program (ILP):
$\min_{\bm{\Psi}\in\{0,1\}^{m\times d_{\Phi}}}\lVert B\hat{\mathbf{K}}_{0}^{\top}-\bm{\Pi}\mathbf{S}\bm{\Psi}^{\top}\rVert_{1}\quad\text{s.t.}\leavevmode\nobreak\ \forall i,b:\sum_{\ell\in L^{(b)}}\psi_{i\ell}=1,\quad\forall i:\leavevmode\nobreak\ \bigcap_{b\in[B]}R_{i}^{(b)}\neq\emptyset.$ (3)
@@ 4.2 Split Relabeling
This strategy may fare poorly if no axis-aligned split in $\mathcal{Z}$ approximates the target split in $\mathcal{X}$. In principle, we could use any binary decision procedure to route samples at internal nodes. For example, using logistic regression, we could create a more complex partition of $\mathcal{Z}$ into convex polytopes. Of course, this increase in expressive flexibility comes at a cost in complexity. The use of synthetic data allows us to choose the effective sample size for learning splits, which is especially advantageous in deep trees, where few training points are available as depth increases.
@@ 4.3 Nearest Neighbors
Neighbors are weighted in inverse proportion to their diffusion distance from the target $\mathbf{z}_{0}$, producing the weight function $w:\mathcal{Z}\mapsto\Delta^{k-1}$. Let $K\subset[n]$ denote the indices of the selected points and $\{\tilde{\mathbf{x}}_{i}\}_{i\in K}$ the corresponding synthetic inputs. Then the $j$th entry of the decoded vector $\hat{\mathbf{x}}_{0}$ is given by $\hat{x}_{0j}=\sum_{i\in K}\leavevmode\nobreak\ w_{i}(\mathbf{z}_{0})\leavevmode\nobreak\ \tilde{x}_{ij}$. For categorical features, we take the most likely label across all neighbors, weighted by $w$, with ties broken randomly. The $k$-NN decoder comes with similar asymptotic guarantees to the previous methods, without assuming access to a leaf assignment oracle.
@@
5 Experiments
In this section, we present a range of experimental results on visualization, reconstruction and denoising. Further details on all experiments can be found in Appx. B, along with additional results. Code for reproducing all figures and tables is included in the supplement.
@@ 5 Experiments
Next, we compare RFAE’s compression-distortion trade-off against two state-of-the-art neural architectures for autoencoding tabular data (TVAE and TTVAE), along with standard and variational autoencoders (AE and VAE, respectively) that are not optimized for this task. Although there are some other notable deep learning algorithms designed for tabular data (e.g., CTGAN [97], TabSyn [100], and TabPFN [51]), these do not come with inbuilt methods for decoding back to the input space at variable compression ratios. We also do not include eForest [30], another RF based autoencoder, because it only works with a fixed $d_{\mathcal{Z}}=2d_{\mathcal{X}}$ and is not capable of compression. For RFAE, we use the unsupervised ARF algorithm [94] with $500$ trees and set $k=20$ for decoding.
In standard AEs, reconstruction error is generally estimated via $L_{2}$ loss. This is not sensible with a mix of continuous and categorical data, so we create a combined measure that evaluates distortion on continuous variables via $1-R^{2}$ (i.e., the proportion of variance unexplained) and categorical variables via classification error. Since both measures are on the unit interval, so too is their average across all $d_{\mathcal{X}}$ features. We plot this measure across a range of compression factors (i.e., inverse compression ratios $d_{\mathcal{Z}}/d_{\mathcal{X}}$) in 20 benchmark tabular datasets (see Fig. 4). For more details on these datasets, see Appx. B.1, Table 1. We evaluate performance over ten bootstrap samples at each compression factor, testing on the randomly excluded out-of-bag data. We find that RFAE is competitive in all settings, and has best average performance in 12 out of 20 datasets (see Appx. B.1, Table 2).

### v2

@@ 2 Background
Our starting point is the well established connection between RFs and kernel methods [16, 28, 83]. The basic insight is that classification and regression trees (CART) [18], which serve as basis functions for many popular ensemble methods, are a kind of adaptive nearest neighbors algorithm [66]. At the root of the tree, all samples are connected. Each split severs the link between one subset of the data and another (i.e., samples routed to left vs. right child nodes), resulting in a gradually sparser graph as depth increases. At completion, a sample’s “neighbors” are just those datapoints that are routed to the same leaf. Given some feature space $\mathcal{X}\subset\mathbb{R}^{d_{\mathcal{X}}}$, the implicit kernel of tree $b$, $k^{(b)}:\mathcal{X}\times\mathcal{X}\mapsto\{0,1\}$, is an indicator function that evaluates to 11 for all and only neighboring sample pairs. [^1] This ignores certain subtleties that arise when trees are grown on bootstrap samples, in which case $k^{(b)}$ may occasionally evaluate to larger integers. For present purposes, we presume that trees are grown on data subsamples; see Appx. A. This base kernel can be used to define different ensemble kernels. For instance, taking an average over $B>1$ trees, we get a kernel with a simple interpretation as the proportion of trees in which two samples colocate: $k^{S}(\mathbf{x},\mathbf{x}^{\prime})=B^{-1}\sum_{b=1}^{B}k^{(b)}(\mathbf{x},\mathbf{x}^{\prime})$. We call this the Scornet kernel after one of its noted proponents [83], who showed that $k^{S}$ is provably close in expectation to the RF similarity function:
@@
where $i\in[n]:=\{1,\dots,n\}$ indexes the training samples. This represents the average of normalized tree kernels, and fully encodes the information learned by the RF $f_{n}$ via the identity:
$f_{n}(\mathbf{x})=\sum_{i=1}^{n}k^{RF}_{n}(\mathbf{x},\mathbf{x}_{i})\penalty 10000\ y_{i},$ (2)
@@
Previous authors have explored feature engineering with RFs. Shi and Horvath [85] perform multidimensional scaling on a dissimilarity matrix extracted from supervised and unsupervised forests. However, they do not explore the connections between this approach and kernel methods, nor do they propose any strategy for decoding latent representations. More heuristic approaches involve running PCA on a weighted matrix of all forest nodes (not just leaves), a method that works well in some experiments but comes with no formal guarantees [77].
Several generative algorithms based on RFs have been proposed, building on the probabilistic circuit literature [27, 99]. These do not involve any explicit encoding step, although the models they train could be passed through our pipeline to extract latent embeddings (see Sect. 5). Existing methods for sum-product network encoding could in principle be applied to an RF following compilation into a corresponding circuit [95]. However, these can often increase rather than decrease dimensionality, and do not come with any associated decoding procedure.
Recent work has explored autoencoding for tree- and forest-based models. Carreira-Perpiñán and Gazizov [20] introduce a sparse oblique tree where each node performs a local linear reconstruction via PCA, yielding a tree-structured autoencoder without use of kernels or ensembles. Aumon et al. [4] propose an RF-informed neural autoencoder, which extracts forest-induced sample similarities to inject supervised signals into network embeddings for data visualization. By contrast, we learn latent representations directly from the RF and implement explicit decoding rules for both supervised and unsupervised settings without an auxiliary model.
Perhaps the most similar method to ours, in motivation if not in execution, is Feng and Zhou’s encoder forest (eForest) [34]. This algorithm maps each point to a hyperrectangle defined by the intersection of all leaves to which the sample is routed. Decoding is then achieved by taking some representative value for each feature in the subregion (e.g., the median). Notably, this approach does not include any dimensionality reduction. On the contrary, the embedding space requires minima and maxima for all input variables, resulting in a representation with double the number of features as the inputs. Working from the conjecture that optimal prediction is equivalent to optimal compression [78, 58, 49], we aim to represent the information learned by the RF in relatively few dimensions.
@@ 3 Encoding
As a preliminary motivation, we prove that the RF similarity function is a proper kernel with several notable properties. The following definitions are standard in the literature. Let $\mathcal{X}$ be a compact metric space, and let $C(\mathcal{X})$ be the set of all real-valued continuous functions on $\mathcal{X}$.
@@
A symmetric function $k:\mathcal{X}\times\mathcal{X}\mapsto\mathbb{R}$ is positive semidefinite (PSD) if, for all $\mathbf{x}_{1},\dots,\mathbf{x}_{n}\in\mathcal{X}$, $n\in\mathbb{N}$, and $c_{i},c_{j}\in\mathbb{R}$, we have: $\sum_{i,j}^{n}c_{i}\penalty 10000\ c_{j}\penalty 10000\ k(\mathbf{x}_{i},\mathbf{x}_{j})\geq 0$.
@@
The bounded measurable kernel $k$ is characteristic if the function $\mu\mapsto\int_{\mathcal{X}}k(\cdot,\mathbf{x})\penalty 10000\ d\mu(\mathbf{x})$ is injective, where $\mu$ is a Borel probability measure on $\mathcal{X}$.
@@
$\lim_{n\rightarrow\infty}\penalty 10000\ P\big(\lVert f^{*}-f_{n}\rVert_{\infty}\geq\epsilon\big)=0.$
@@
The literature on kernel methods is largely focused on fixed kernels such as the radial basis function (RBF). Among adaptive partitioning alternatives, the Scornet kernel been studied in some detail [28, 83, 76], as has the Mondrian kernel [64, 73, 21]. However, to the best of our knowledge, we are the first to establish these properties for the RF kernel. Thm. 3.4 confirms that $k_{n}^{RF}$ is flexible, informative, and generally “well-behaved” in ways that will prove helpful for autoencoding.
@@
$\min_{\mathbf{V}\in\mathbb{R}^{n\times d_{\mathcal{Z}}}}\sum_{i,j}k_{ij}\penalty 10000\ \lVert\mathbf{v}_{i}-\mathbf{v}_{j}\rVert_{2}\quad\text{s.t.}\penalty 10000\ \mathbf{V}^{\top}\mathbf{V}=\mathbf{I},$
@@ 4 Decoding
$\lim_{n\rightarrow\infty}\penalty 10000\ P\Big(\lVert\mathbf{x}-h_{n}\big(g^{*}(\mathbf{x})\big)\rVert_{\infty}\geq\epsilon\Big)=0.$
@@
Let $f_{n}$ be a RF trained on $\{\mathbf{x}_{i},y_{i}\}_{i=1}^{n}\overset{\text{i.i.d.}}{\sim}P_{XY}$. Let $h_{n}^{*}:\mathcal{Z}\mapsto\mathcal{X}$ be a decoder that (i) maps latent vectors to leaves in $f_{n}$ via the oracle $\psi^{*}_{f_{n},g}$; then (ii) reconstructs data by sampling uniformly from the intersection of assigned leaves for each sample. Then, under the assumptions of Thm. 3.4, the sequence $\{h^{*}_{n}\}$ is universally consistent.
@@ 4.1 Constrained Optimization
Now we must identify the most likely leaf assignments for each of our $m$ (unseen) test samples $\mathbf{X}_{0}\in\mathbb{R}^{m\times d_{\mathcal{X}}}$, given their latent representation $\mathbf{Z}_{0}$. Call this target matrix $\bm{\Psi}\in\{0,1\}^{m\times d_{\Phi}}$. To estimate it, we start with the binary matrix of leaf assignments for training samples $\bm{\Pi}\in\{0,1\}^{n\times d_{\Phi}}$. Exploiting the inner product definition of a kernel, observe that our original (training) adjacency matrix $\mathbf{K}$ satisfies $B\mathbf{K}=\bm{\Phi}\bm{\Phi}^{\top}=\bm{\Pi}\mathbf{S}\bm{\Pi}^{\top}$, where $\bm{\Phi}\in[0,1]^{n\times d_{\Phi}}$ is the RKHS representation of $\mathbf{X}$, and $\mathbf{S}=\text{diag}(\mathbf{s})$. We partition $d_{\Phi}]$ into $B$ subsets $L^{(b)}$ that index the leaves belonging to tree $b$. Recall that each tree $b$ partitions the feature space $\mathcal{X}$ into $L^{(b)}$ hyperrectangular subregions $\mathcal{X}_{\ell}^{(b)}\subset\mathcal{X}$, one for each leaf $\ell\in[L^{(b)}]$. Let $R_{i}^{(b)}\in\{\mathcal{X}_{1}^{(b)},\dots,\mathcal{X}_{L^{(b)}}^{(b)}\}$ denote the region to which sample $i$ is routed in tree $b$. Then leaf assignments for test samples can be calculated by solving the following integer linear program (ILP):
$\min_{\bm{\Psi}\in\{0,1\}^{m\times d_{\Phi}}}\lVert B\hat{\mathbf{K}}_{0}^{\top}-\bm{\Pi}\mathbf{S}\bm{\Psi}^{\top}\rVert_{1}\quad\text{s.t.}\penalty 10000\ \forall i,b:\sum_{\ell\in L^{(b)}}\psi_{i\ell}=1,\quad\forall i:\penalty 10000\ \bigcap_{b\in[B]}R_{i}^{(b)}\neq\emptyset.$ (3)
@@ 4.2 Split Relabeling
This strategy may fare poorly if no axis-aligned split in $\mathcal{Z}$ approximates the target split in $\mathcal{X}$. In principle, we could use any binary decision procedure to route samples at internal nodes. For example, using logistic regression, we could create a more complex partition of $\mathcal{Z}$ into convex polytopes. Of course, this increase in expressive flexibility comes at a cost in space and time complexity. The use of synthetic data allows us to choose the effective sample size for learning splits, which is especially advantageous in deep trees, where few training points are available as depth increases.
@@ 4.3 Nearest Neighbors
Neighbors are weighted in inverse proportion to their diffusion distance from the target $\mathbf{z}_{0}$, producing the weight function $w:\mathcal{Z}\mapsto\Delta^{k-1}$. Let $K\subset[n]$ denote the indices of the selected points and $\{\tilde{\mathbf{x}}_{i}\}_{i\in K}$ the corresponding synthetic inputs. Then the $j$th entry of the decoded vector $\hat{\mathbf{x}}_{0}$ is given by $\hat{x}_{0j}=\sum_{i\in K}\penalty 10000\ w_{i}(\mathbf{z}_{0})\penalty 10000\ \tilde{x}_{ij}$. For categorical features, we take the most likely label across all neighbors, weighted by $w$, with ties broken randomly. The $k$-NN decoder comes with similar asymptotic guarantees to the previous methods, without assuming access to a leaf assignment oracle.
@@
5 Experiments
In this section, we present a range of experimental results on visualization, reconstruction and denoising. Further details on all experiments can be found in Appx. B, along with additional results. Code for reproducing all results is available online. [^4] https://github.com/bips-hb/RFAE.
@@ 5 Experiments
Table 1: Mean distortion (with standard errors) for each method and dataset, across all $d_{\mathcal{Z}}$ values used and all runs. Best result per row is bolded.
Dataset RFAE TVAE TTVAE AE VAE
abalone 0.167 (0.002) 0.309 (0.005) 0.260 (0.003) 0.230 (0.025) 0.211 (0.006)
adult 0.326 (0.007) 0.158 (0.005) 0.195 (0.007) 0.401 (0.003) 0.391 (0.004)
banknote 0.100 (0.012) 0.312 (0.013) 0.276 (0.023) 0.724 (0.023) 0.771 (0.013)
bc 0.333 (0.003) 0.564 (0.003) 0.359 (0.005) 0.287 (0.008) 0.578 (0.003)
car 0.320 (0.011) 0.195 (0.014) 0.107 (0.015) 0.349 (0.012) 0.313 (0.011)
churn 0.352 (0.012) 0.603 (0.011) 0.422 (0.014) 0.861 (0.005) 0.731 (0.006)
credit 0.315 (0.004) 0.450 (0.005) 0.375 (0.011) 0.450 (0.005) 0.456 (0.004)
diabetes 0.479 (0.016) 0.726 (0.007) 0.643 (0.014) 0.799 (0.011) 0.895 (0.004)
dry_bean 0.137 (0.002) 0.273 (0.002) 0.303 (0.008) 0.083 (0.014) 0.206 (0.001)
hd 0.432 (0.008) 0.582 (0.003) 0.605 (0.006) 0.892 (0.003) 0.916 (0.002)
king 0.308 (0.008) 0.352 (0.006) 0.348 (0.008) 0.377 (0.011) 0.518 (0.004)
marketing 0.292 (0.009) 0.304 (0.005) 0.259 (0.011) 0.357 (0.007) 0.372 (0.004)
mushroom 0.083 (0.001) 0.093 (0.003) 0.011 (0.003) 0.055 (0.004) 0.035 (0.004)
obesity 0.227 (0.008) 0.354 (0.004) 0.299 (0.008) 0.306 (0.009) 0.358 (0.003)
plpn 0.176 (0.006) 0.282 (0.006) 0.224 (0.011) 0.384 (0.013) 0.410 (0.009)
spambase 0.558 (0.005) 0.825 (0.002) 0.807 (0.003) 0.446 (0.010) 0.784 (0.001)
student 0.371 (0.002) 0.424 (0.001) 0.426 (0.004) 0.536 (0.003) 0.551 (0.002)
telco 0.177 (0.003) 0.155 (0.003) 0.091 (0.007) 0.128 (0.005) 0.130 (0.005)
wq 0.240 (0.005) 0.691 (0.008) 0.759 (0.006) 0.467 (0.019) 0.708 (0.004)
Average Rank 1.80 3.38 2.45 3.27 4.10
@@
Next, we compare RFAE’s compression-distortion trade-off against two state-of-the-art neural architectures for autoencoding tabular data (TVAE and TTVAE), along with standard and variational autoencoders (AE and VAE, respectively) that are not optimized for this task. Although there are some other notable deep learning algorithms designed for tabular data (e.g., CTGAN [103], TabSyn [106], and TabPFN [56]), these do not come with inbuilt methods for decoding back to the input space at variable compression ratios. We also do not include eForest [34], another RF based autoencoder, because it only works with a fixed $d_{\mathcal{Z}}=2d_{\mathcal{X}}$ and is not capable of compression. For RFAE, we use the unsupervised ARF algorithm [99] with 500500 trees and set $k=20$ for decoding.
In standard AEs, reconstruction error is generally estimated via $L_{2}$ loss. This is not sensible with a mix of continuous and categorical data, so we create a combined measure that evaluates distortion on continuous variables via $1-R^{2}$ (i.e., the proportion of variance unexplained) and categorical variables via classification error. Since both measures are on the unit interval, so too is their average across all $d_{\mathcal{X}}$ features.
We measure this distortion across a range of compression factors (i.e., inverse compression ratios $d_{\mathcal{Z}}/d_{\mathcal{X}}$ in 20 benchmark tabular datasets. For more details on these datasets, see Appx. B.1, Table 2. We evaluate performance over ten bootstrap samples at each compression factor, testing on the randomly excluded out-of-bag data. We find that RFAE is competitive in all settings, and has best average performance in 12 out of 20 datasets (see Table 1), for an average rank of 1.80. For a more granular view of performance on each dataset, see Appx. B.2, Fig. 5.